5 points

Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
Direct materials 3 feet $ 5 per foot $ 15
Direct labor ? hours ? per hour ?

During March, the company purchased direct materials at a cost of $57,090, all of which were used in the production of 3,400 units of product. In addition, 5,200 direct labor-hours were worked on the product during the month. The cost of this labor time was $54,600. The following variances have been computed for the month:

Materials quantity variance $ 900 U
Labor spending variance $ 3,600

U

Labor efficiency variance $ 1,000

U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.
Answer #1

Solution 1a (Sharp Company):

Material quantity variance = $900 U

(SQ - AQ)*SP = -$900

(3400*3 - AQ) * $5 = - $900

Actual quantity = 10380 Foot

Actual cost of purchase = $57,090

actual cost per foot = $57090 / 10380 = $5.50 per foot

Solution 1b:

Material price variance = (SP - AP) * AQ = ($5 - $5.50) * 10380 = $5,190 U

Material spending variance = MPV + MUV = $5190 U + $900 U = $6,090 U

Solution 2a:

Actual rate of labor per hour = $54600 / 5200 = $10.50 per hour

Labor rate variance = Labor spending variance - Labor efficiency variance = $3,600 U - $1,000 U = $2,600 U

(SR - AR) * AH = - $2600

(SR - $10.50) * 5200 = - $2600

SR = $10 per hour

Solution 2b:

Labor efficiency variance = (SH - AH) * SR = (SH - 5200) * $10 = - $1000

Standard hours = 5100 hours

solution 2c:

Standard hours allowed per unit = 5100 / 3400 = 1.50 hour per unit.